# Bibliografia.bib

### v1

@@ 1. Introduction
  • quadratic rational if $[{\mathbb{Q}}(\chi):{\mathbb{Q}}]\leq 2$ for every $\chi\in\operatorname{Irr}(G)$.
@@
  • quasi-rational if $[{\mathbb{Q}}(G):{\mathbb{Q}}]\leq 2$;
@@
RationalInverse semi-rationalQuasi-rational  Prop. 3.13 USR  Cor. 3.12 Semi-rationalQuadratic rationalQuadratic valued Cor. 3.4 Quadratic conjugated
@@ 2. Notation
We use standard notation $C_{n}$ and $D_{n}$ for cyclic and dihedral groups, respectively, of order $n$. Moreover, $\left\langle a\right\rangle_{n}$ represent a cyclic group of order $n$ generated by $a$. Other groups which appear in this paper are the semi-dihedral groups, for $k\geq 3$:
@@
Let $\operatorname{Irr}(G)$ denote the set of irreducible ordinary characters of $G$. If $\chi\in\operatorname{Irr}(G)$, then ${\mathbb{Q}}(\chi)={\mathbb{Q}}(\chi(x):x\in G)$, the field of character values of $\chi$. If $g\in G$, then ${\mathbb{Q}}(g)={\mathbb{Q}}(\psi(g):\psi\in\operatorname{Irr}(G))$ and $T(g):\operatorname{Irr}(G)\to{\mathbb{C}}$ is the map given by $T(g)(\chi)=\chi(g)$. Observe that $T(g)=T(h)$ if and only if $g\sim h$. Let $T(G)$ denote the set of maps $T(g)$ with $g\in G$ and set ${\mathbb{Q}}(G)={\mathbb{Q}}(\chi(g):\chi\in\operatorname{Irr}(G),g\in G)$.
@@
  • $C_{3}$ and $C_{4}$ are inverse semi-rational and quasi-rational but not rational.
  • $D_{16}$ is quasi-rational but not inverse semi-rational.
  • $\left\langle a\right\rangle_{8}\rtimes\left\langle b\right\rangle_{4}$ with $b^{a}=a^{3}$, is inverse semi-rational but not quasi-rational.
  • $\left\langle a\right\rangle_{8}\rtimes\left\langle b\right\rangle_{4}$ with $b^{a}=a^{-1}$, is USR but neither inverse semi-rational nor quasi-rational.
  • SmallGroup(128,417)$=\left\langle a\right\rangle_{8}\rtimes(\left\langle b\right\rangle_{8}\rtimes\left\langle c\right\rangle_{2})$, with $a^{b}=a^{-1}$, $a^{c}=a^{3}$, $b^{c}=b^{-1}$, is semi-rational and quadratic rational but not USR.
  • SmallGroup(32,15)=$\left\langle a,b\mid a^{8}=1,b^{4}=a^{4},a^{b}=a^{3}\right\rangle$ is semi-rational but not quadratic rational.
  • SmallGroup(32,9) $=(\left\langle a\right\rangle_{8}\times\left\langle b\right\rangle_{2})\rtimes\left\langle c\right\rangle_{2}$, with $a^{c}=ab$, $b^{c}=a^{4}b$ is quadratic rational but not semi-rational.
  • SmallGroup(64,22) is quadratic valued but neither semi-rational nor quadratic rational.
@@ 3.1. Basic properties
Combining this with Lemma 3.1 it follows that if $g$ is semi-rational in $G$, then $\mathcal{S}_{g}$ is a coset of $\mathcal{R}_{g}$ and $\mathcal{U}_{G}^{2}\subseteq\mathcal{R}_{g}$. As an intersection of cosets of a group is either empty or a coset modulo the intersection of the corresponding subgroups we obtain the following.
@@ 3.2. Groups inert over a field
Let $K$ be a subfield of ${\mathbb{C}}$. We say that $G$ is inert over $K$ if for every irreducible ${\mathbb{Q}}G$-module $M$, the $KG$-module $K\otimes_{\mathbb{Q}}M$ is irreducible.
@@
By the Witt-Berman Theorem [CurtisReiner, Theorem 42.8], the number of $K$-conjugacy classes of $G$ coincides with the number of isomorphism classes of irreducible $KG$-modules and therefore $G$ is inert over $K$ if and only if $g^{G}_{K}=g^{G}_{\mathbb{Q}}$ for every $g\in G$. Hence (1) and (2) are equivalent.
@@
As the sets $\operatorname{Irr}(G)$ and $T(G)$ are linearly independent over ${\mathbb{C}}$ and invariant under the action of $\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{Q}})$, it follows that there are $\tau_{\chi}$ and $\rho_{g}$ in $\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{Q}})$ such that $\chi^{\sigma}=\chi^{\tau_{\chi}}$ and $T(g)^{\sigma}=T(g)^{\rho_{g}}$.
@@
Finally, we prove that (8) implies (2). Suppose that ${\mathbb{Q}}(\chi)\cap K={\mathbb{Q}}$ for every $\chi\in\operatorname{Irr}(G)$. Let $A$ be the unique simple component of ${\mathbb{Q}}G$ with $\chi(A)\neq 0$. Then ${\bf Z}(A)={\mathbb{Q}}(\chi)$ (see e.g. [JespersdelRioGRG1, Theorem 3.3.1(4)]) and the assumption ${\mathbb{Q}}(\chi)\cap K={\mathbb{Q}}$ implies that $K\otimes_{{\mathbb{Q}}}A$ is a central simple $K(\chi)$-algebra. This implies that ${\mathbb{Q}}G$ and $KG$ have the same number of simple components. By the Witt-Berman Theorem, the number of simple components of $KG$ is the number of $K$-classes of $G$. Thus the number of ${\mathbb{Q}}$-classes and $K$-conjugacy classes of $G$ coincides and hence $g^{G}_{\mathbb{Q}}=g^{G}_{K}$ for every $g\in G$.
@@ 3.3. Uniformly semi-rational groups
    $G$ is quadratic valued and inert over ${\mathbb{K}}_{r}$.
@@
    $G$ is quadratic conjugated and inert over ${\mathbb{K}}_{r}$.
@@
We start by proving (1) implies (2). Suppose that $G$ is $r$-semi-rational. Thus $G$ is quadratic valued. Moreover, for every $g\in G$ we have $g^{G}_{{\mathbb{K}}_{r}}\subseteq g^{G}_{\mathbb{Q}}=g^{G}\cup(g^{r})^{G}\subseteq g^{G}_{{\mathbb{K}}_{r}}$. Therefore, by Proposition 3.6, $G$ is inert over ${\mathbb{K}}_{r}$.
@@
To prove (3) implies (1), we assume that $G$ is quadratic conjugated and inert over ${\mathbb{K}}_{r}$. Let $g\in G$. We have to prove that $g$ is $r$-semi-rational in $G$. On the one hand, $\operatorname{Gal}({\mathbb{Q}}(g)/{\mathbb{Q}})$ is elementary abelian $2$-group, by Proposition 3.3; on the other hand, the restriction map $\left\langle\sigma_{r}\right\rangle=\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{K}}_{r})\to\operatorname{Gal}({\mathbb{Q}}(g)/{\mathbb{Q}})$ is surjective, by Proposition 3.6. Then $[{\mathbb{Q}}(g):{\mathbb{Q}}]\leq 2$, and hence $g$ is semi-rational in $G$ by Lemma 3.1. If $g$ is rational in $G$, then it is $r$-semi-rational in $G$. Otherwise $[{\mathbb{Q}}(g):{\mathbb{Q}}]=2$ and, by Proposition 3.6, ${\mathbb{Q}}(g)\cap{\mathbb{K}}_{r}={\mathbb{Q}}$. Therefore ${\mathbb{Q}}(g)$ is not contained in ${\mathbb{K}}_{r}$ and hence $\chi(g^{r})=\sigma_{r}(\chi(g))\neq\chi(g)$ for some $\chi\in{\rm Irr}(G)$. Thus $g\not\sim g^{r}$, proving that $g$ is $r$-semi-rational in $G$, as desired.
@@
Let $r\in\mathcal{U}_{G}$ with $r^{2}=1$. Then $G$ is $r$-semi-rational if and only if $G$ inert over ${\mathbb{K}}_{r}$.
@@
The necessary condition is consequence of Proposition 3.7. To prove the sufficient condition, we assume that $r^{2}=1$ and $G$ is inert over ${\mathbb{K}}_{r}$. Then the order of $\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{K}}_{r})$ is at most $2$ and, by Proposition 3.6, we get that $\operatorname{Gal}({\mathbb{Q}}(\chi)/{\mathbb{Q}})$ is also of order at most $2$ for every $\chi\in\operatorname{Irr}(G)$. Then $G$ is quadratic valued and, by Proposition 3.7, $G$ is $r$-semi-rational.
@@
The dihedral group $D_{32}$ of order $32$ is inert over ${\mathbb{K}}_{3}={\mathbb{Q}}(\sqrt{-2})$, but it is not USR because $G$ is not quadratic conjugated, as the conjugacy class of an element $g$ of order $16$ in $D_{32}$ does not contain $g^{9}$.
@@
The group $G$ is inverse semi-rational if and only if $G$ is inert over $\mathbb{R}$.
@@
    $G$ is quadratic valued and inert over $K$ for some subfield $K$ of ${\mathbb{Q}}_{G}$ such that ${\mathbb{Q}}_{G}/K$ is a cyclic extension.
@@
    $G$ is quadratic conjugated and inert over for some subfield $K$ of ${\mathbb{Q}}_{G}$ such that ${\mathbb{Q}}_{G}/K$ is a cyclic extension.
A now follows from Proposition 3.6, Proposition 3.7 and Corollary 3.11. Thought, not all the semi-rational groups are quadratic rational, from Corollary 3.11 we obtain the following corollary.
@@
If $G$ is USR, then it is quadratic rational.
@@
Suppose that $G$ is USR. By Corollary 3.11, $G$ is quadratic conjugated and there exists a subfield $K$ of ${\mathbb{Q}}_{G}$ such that $G$ is inert over $K$ and ${\mathbb{Q}}_{G}/K$ is a cyclic extension. Then for every $\chi\in{\rm Irr}(G)$, $\operatorname{Gal}({\mathbb{Q}}(\chi)/{\mathbb{Q}})$ is elementary abelian $2$-group, by Proposition 3.3, and the restriction map $\operatorname{Gal}({\mathbb{Q}}_{G}/K)\to\operatorname{Gal}({\mathbb{Q}}(\chi)/{\mathbb{Q}})$ is surjective, by Proposition 3.6. Therefore $\operatorname{Gal}({\mathbb{Q}}(\chi)/{\mathbb{Q}})$ is cyclic of order at most $2$, implying that $[{\mathbb{Q}}(\chi):{\mathbb{Q}}]\leq 2$. This proves that $G$ is quadratic rational.
3.4. Quasi-rational groups and direct products
In this subsection we obtain two characterizations of quasi-rational groups. In the way, we characterize when the direct product of two groups is USR. Our first characterization follows by elementary arguments.
@@ 3.4. Quasi-rational groups and direct products
    ${\bf Z}({\mathbb{Q}}G)\cong{\mathbb{Q}}^{n}\times K^{m}$ for an extension $K$ of ${\mathbb{Q}}$ of degree at most $2$ for some non-negative integers $n,m$.
@@
Moreover, if $K$ is a subfield of ${\mathbb{Q}}_{H}$, then the following diagram is commutative, where Res stands for a restriction map.
@@
    $G$ and $H$ are quasi-rational and ${\mathbb{Q}}(G)={\mathbb{Q}}(H)$.
In that case $\mathcal{R}_{G\times H}=\pi^{-1}_{G\times H,G}(\mathcal{R}_{G})\cap\pi^{-1}_{G\times H,H}(\mathcal{R}_{H})$ and $\mathcal{S}_{G\times H}=\pi^{-1}_{G\times H,G}(\mathcal{S}_{G})\cap\pi^{-1}_{G\times H,H}(\mathcal{S}_{H})$. Moreover, if (2) holds, then $G\times H$ is quasi-rational and ${\mathbb{Q}}(G\times H)={\mathbb{Q}}(G)$.
@@
Now suppose that $G$ and $H$ are quasi-rational and ${\mathbb{Q}}(G)={\mathbb{Q}}(H)=K$. By the previous case, we may assume that $K\neq{\mathbb{Q}}$, and we have to show that $G\times H$ is quasi-rational, $\mathcal{R}_{G\times H}=\pi^{-1}_{G}(\mathcal{R}_{G})=\pi_{H}^{-1}(\mathcal{R}_{H})$ and $\mathcal{S}_{G\times H}=\pi^{-1}_{G}(\mathcal{S}_{G})=\pi^{-1}_{H}(\mathcal{S}_{H})$. By Proposition 3.13, ${\bf Z}({\mathbb{Q}}G)$ is isomorphic to ${\mathbb{Q}}^{n_{1}}\times K^{m_{1}}$ and ${\bf Z}({\mathbb{Q}}H)$ is isomorphic to ${\mathbb{Q}}^{n_{2}}\times K^{m_{2}}$ for some non-negative integers $n_{1},n_{2},m_{1}$ and $m_{2}$. Then ${\bf Z}({\mathbb{Q}}(G\times H))\cong{\bf Z}({\mathbb{Q}}G)\otimes_{\mathbb{Q}}{\bf Z}({\mathbb{Q}}H)\cong{\mathbb{Q}}^{n}\times K^{m}$ where $n=n_{1}n_{2}$ and $m=n_{1}m_{2}+n_{2}m_{1}+2m_{1}m_{2}$. So, by Proposition 3.13, the group $G\times H$ is quasi-rational and $K={\mathbb{Q}}(G\times H)$. By equations (2.1) and (3.2), we have that $\mathcal{R}_{G\times H}=\sigma_{G\times H}^{-1}(\operatorname{Gal}({\mathbb{Q}}_{G\times H}/K))=\pi^{-1}_{G}(\sigma_{G}^{-1}(\operatorname{Gal}({\mathbb{Q}}_{G}/K)))=\pi^{-1}_{G}(\mathcal{R}_{G})$. Moreover, by Proposition 3.13,
@@
Let $G$ be a USR group which is not quasi-rational. If $G=G_{1}\times G_{2}$, then either $G_{1}$ or $G_{2}$ is rational.
We conclude the section with the second characterization of quasi-rational groups.

### v2

@@ 1. Introduction
  • quadratic if $[{\mathbb{Q}}(G):{\mathbb{Q}}]\leq 2$;
  • character quadratic if $[{\mathbb{Q}}(\chi):{\mathbb{Q}}]\leq 2$ for every $\chi\in\operatorname{Irr}(G)$. In [tentquadraticrat], it is named quadratic rational.
@@
RationalInverse semi-rationalQuadratic  Prop. 3.13 USR  Cor. 3.12 Semi-rationalCharacter quadraticQuadratic valued Cor. 3.4 Quadratic conjugated
@@ 2. Notation
We use standard notation $C_{n}$ and $D_{n}$ for cyclic and dihedral groups, respectively, of order $n$. Moreover, $\left\langle a\right\rangle_{n}$ represents a cyclic group of order $n$ generated by $a$. Other groups which appear in this paper are the semi-dihedral groups, for $k\geq 3$:
@@
If $g\in G$, then $T(g):\operatorname{Irr}(G)\to{\mathbb{C}}$ is the map given by $T(g)(\chi)=\chi(g)$. Observe that $T(g)=T(h)$ if and only if $g\sim h$. Let $T(G)$ denote the set of maps $T(g)$ with $g\in G$.
@@
  • $C_{3}$ and $C_{4}$ are inverse semi-rational and quadratic but not rational.
  • $D_{16}$ is quadratic but not inverse semi-rational.
  • $\left\langle a\right\rangle_{8}\rtimes\left\langle b\right\rangle_{4}$ with $b^{a}=a^{3}$, is inverse semi-rational but not quadratic.
  • $\left\langle a\right\rangle_{8}\rtimes\left\langle b\right\rangle_{4}$ with $b^{a}=a^{-1}$, is USR but neither inverse semi-rational nor quadratic.
  • SmallGroup(128,417)$=\left\langle a\right\rangle_{8}\rtimes(\left\langle b\right\rangle_{8}\rtimes\left\langle c\right\rangle_{2})$, with $a^{b}=a^{-1}$, $a^{c}=a^{3}$, $b^{c}=b^{-1}$, is semi-rational and character quadratic but not USR.
  • SmallGroup(32,15)=$\left\langle a,b\mid a^{8}=1,b^{4}=a^{4},a^{b}=a^{3}\right\rangle$ is semi-rational but not character quadratic.
  • SmallGroup(32,9) $=(\left\langle a\right\rangle_{8}\times\left\langle b\right\rangle_{2})\rtimes\left\langle c\right\rangle_{2}$, with $a^{c}=ab$, $b^{c}=a^{4}b$ is character quadratic but not semi-rational.
  • SmallGroup(64,22) is quadratic valued but neither semi-rational nor character quadratic.
@@ 3.1. Basic properties
Combining this with Lemma 3.1 it follows that if $g$ is semi-rational in $G$, then $\mathcal{S}_{g}$ is a coset of $\mathcal{R}_{g}$ in $\mathcal{U}_{G}$ and $\mathcal{U}_{G}^{2}\subseteq\mathcal{R}_{g}$. As an intersection of cosets of a group is either empty or a coset modulo the intersection of the corresponding subgroups we obtain the following.
@@ 3.2. Stillness
Recall that every semisimple module is the direct sum of its homogeneous components. We say that a semisimple module is homogeneous if it has exactly one non-zero homogeneous component. Two homogeneous modules are equivalent if they have isomorphic simple direct summands. So the number of isomorphism classes of simple modules of a ring equals the number of equivalence classes of semisimple homogenous modules. Furthermore, if $A$ is $K$-algebra, $F$ is a field extension of $K$, $H$ is a homogeneous $A$-module and $S$ a simple direct summand of $H$, then $F\otimes_{K}H$ and $F\otimes_{K}S$ have the same number of homogeneous components.
@@
Let $K$ be a subfield of ${\mathbb{C}}$. We say that $G$ is $K$-still if for every homogeneous ${\mathbb{Q}}G$-module $H$, the $KG$-module $K\otimes_{\mathbb{Q}}H$ is homogeneous.
Let $K$ be a field and let $A$ be a semisimple $K$-algebra. Then $A=\oplus_{i=1}^{n}A_{i}$ with each $A_{i}$ a simple $K$-algebra. Then $A_{1},\dots,A_{n}$ is a set of representatives of the equivalence classes of homogeneous $A$-modules. Let now $F$ be a field extension of $K$ and $B=F\otimes_{K}A$. Then $F\otimes_{K}A_{i}=\oplus_{j=1}^{k_{i}}B_{i,j}$ with each $B_{i,j}$ a simple algebra and hence $B=\oplus_{i=1}^{n}\oplus_{j=1}^{k_{i}}B_{i,j}$. Thus the $B_{i,j}$ form a set of representatives of the equivalence classes of homogeneous $B$-modules. Therefore, the number of equivalence classes of homogeneous $B$-modules is $\sum_{i=1}^{n}k_{i}$. This implies that if $K$ is a subfield of ${\mathbb{C}}$, then $G$ is $K$-still if and only if ${\mathbb{Q}}G$ and $KG$ have the same number of equivalence classes of homogeneous modules, equivalently the number of ${\mathbb{Q}}$-characters equals the number of $K$-characters.
@@
By the Witt-Berman Theorem [CurtisReiner, Theorem 42.8], the number of $K$-conjugacy classes of $G$ coincides with the number of isomorphism classes of irreducible $KG$-modules, which is the same as the number of equivalence classes of homogeneous $KG$-modules, and therefore $G$ is $K$-still if and only if $g^{G}_{K}=g^{G}_{\mathbb{Q}}$ for every $g\in G$. Hence (1) and (2) are equivalent.
@@
As the sets $\operatorname{Irr}(G)$ and $T(G)$ are linearly independent over ${\mathbb{C}}$ and invariant under the action of $\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{Q}})$, it follows that there are $\tau_{\chi}$ and $\rho_{g}$ in $\operatorname{Gal}({\mathbb{Q}}_{G}/K)$ such that $\chi^{\sigma}=\chi^{\tau_{\chi}}$ and $T(g)^{\sigma}=T(g)^{\rho_{g}}$.
@@
Finally, we prove that (8) implies (2). Suppose that ${\mathbb{Q}}(\chi)\cap K={\mathbb{Q}}$ for every $\chi\in\operatorname{Irr}(G)$. Let $\chi\in\operatorname{Irr}(G)$. Then $[{\mathbb{Q}}(\chi):{\mathbb{Q}}]=[K(\chi):K]$. Let $A$ be the unique simple component of ${\mathbb{Q}}G$ with $\chi(A)\neq 0$. By items (4) and (5) in [JespersdelRioGRG1, Theorem 3.3.1], we get that ${\bf Z}(A)={\mathbb{Q}}(\chi)$ and $K\otimes_{{\mathbb{Q}}}A$ is a central simple $K(\chi)$-algebra. This implies that ${\mathbb{Q}}G$ and $KG$ have the same number of simple components. By the Witt-Berman Theorem, the number of simple components of $KG$ is the number of $K$-classes of $G$. Thus the number of ${\mathbb{Q}}$-classes and $K$-conjugacy classes of $G$ coincides and hence $g^{G}_{\mathbb{Q}}=g^{G}_{K}$ for every $g\in G$.
@@ 3.3. Uniformly semi-rational groups
    $G$ is quadratic valued and ${\mathbb{K}}_{r}$-still.
@@
    $G$ is quadratic conjugated and ${\mathbb{K}}_{r}$-still.
@@
We start by proving (1) implies (2). Suppose that $G$ is $r$-semi-rational. Thus $G$ is quadratic valued. Moreover, for every $g\in G$ we have $g^{G}_{{\mathbb{K}}_{r}}\subseteq g^{G}_{\mathbb{Q}}=g^{G}\cup(g^{r})^{G}\subseteq g^{G}_{{\mathbb{K}}_{r}}$. Therefore, by Proposition 3.6, $G$ is ${\mathbb{K}}_{r}$-still.
@@
To prove (3) implies (1), we assume that $G$ is quadratic conjugated and ${\mathbb{K}}_{r}$-still. Let $g\in G$. We have to prove that $g$ is $r$-semi-rational in $G$. On the one hand, $\operatorname{Gal}({\mathbb{Q}}(g)/{\mathbb{Q}})$ is elementary abelian $2$-group, by Proposition 3.3; on the other hand, the restriction map $\left\langle\sigma_{r}\right\rangle=\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{K}}_{r})\to\operatorname{Gal}({\mathbb{Q}}(g)/{\mathbb{Q}})$ is surjective, by Proposition 3.6. Then $[{\mathbb{Q}}(g):{\mathbb{Q}}]\leq 2$, and hence $g$ is semi-rational in $G$ by Lemma 3.1. If $g$ is rational in $G$, then it is $r$-semi-rational in $G$. Otherwise $[{\mathbb{Q}}(g):{\mathbb{Q}}]=2$ and, by Proposition 3.6, ${\mathbb{Q}}(g)\cap{\mathbb{K}}_{r}={\mathbb{Q}}$. Therefore ${\mathbb{Q}}(g)$ is not contained in ${\mathbb{K}}_{r}$ and hence $\chi(g^{r})=\sigma_{r}(\chi(g))\neq\chi(g)$ for some $\chi\in{\rm Irr}(G)$. Thus $g\not\sim g^{r}$, proving that $g$ is $r$-semi-rational in $G$, as desired.
@@
Let $r\in\mathcal{U}_{G}$ with $r^{2}=1$. Then $G$ is $r$-semi-rational if and only if $G$ is ${\mathbb{K}}_{r}$-still.
@@
The necessary condition is consequence of Proposition 3.7. To prove the sufficient condition, we assume that $r^{2}=1$ and $G$ is ${\mathbb{K}}_{r}$-still. Then the order of $\operatorname{Gal}({\mathbb{Q}}_{G}/{\mathbb{K}}_{r})$ is at most $2$ and, by Proposition 3.6, we get that $\operatorname{Gal}({\mathbb{Q}}(\chi)/{\mathbb{Q}})$ is also of order at most $2$ for every $\chi\in\operatorname{Irr}(G)$. Then $G$ is quadratic valued and, by Proposition 3.7, $G$ is $r$-semi-rational.
@@
The dihedral group $D_{32}$ of order $32$ is ${\mathbb{K}}_{3}$-still (for this group ${\mathbb{K}}_{3}={\mathbb{Q}}(\sqrt{-2})$), but it is not USR because $G$ is not quadratic conjugated, as the conjugacy class of an element $g$ of order $16$ in $D_{32}$ does not contain $g^{9}$.
@@
The group $G$ is inverse semi-rational if and only if $G$ is $\mathbb{R}$-still.
@@
    $G$ is quadratic valued and $K$-still for some subfield $K$ of ${\mathbb{Q}}_{G}$ such that ${\mathbb{Q}}_{G}/K$ is a cyclic extension.
@@
    $G$ is quadratic conjugated and $K$-still for some subfield $K$ of ${\mathbb{Q}}_{G}$ such that ${\mathbb{Q}}_{G}/K$ is a cyclic extension.
A now follows from Proposition 3.6 and Corollary 3.11. Thought, not all the semi-rational groups are character quadratic, from Corollary 3.11 we obtain the following corollary.
@@
If $G$ is USR, then it is character quadratic.
@@
Suppose that $G$ is USR. By Corollary 3.11, $G$ is quadratic conjugated and there exists a subfield $K$ of ${\mathbb{Q}}_{G}$ such that $G$ is $K$-still and ${\mathbb{Q}}_{G}/K$ is a cyclic extension. Then for every $\chi\in{\rm Irr}(G)$, $\operatorname{Gal}({\mathbb{Q}}(\chi)/{\mathbb{Q}})$ is elementary abelian $2$-group, by Proposition 3.3, and the restriction map $\operatorname{Gal}({\mathbb{Q}}_{G}/K)\to\operatorname{Gal}({\mathbb{Q}}(\chi)/{\mathbb{Q}})$ is surjective, by Proposition 3.6. Therefore $\operatorname{Gal}({\mathbb{Q}}(\chi)/{\mathbb{Q}})$ is cyclic of order at most $2$, implying that $[{\mathbb{Q}}(\chi):{\mathbb{Q}}]\leq 2$. This proves that $G$ is character quadratic.
3.4. Quadratic groups and direct products
In this subsection we obtain two characterizations of quadratic groups. In the way, we characterize when the direct product of two groups is USR. Our first characterization follows by elementary arguments.
@@ 3.4. Quadratic groups and direct products
    ${\bf Z}({\mathbb{Q}}G)\cong{\mathbb{Q}}^{n}\times K^{m}$ for an extension $K$ of ${\mathbb{Q}}$ of degree at most $2$ for some non-negative integer $n$ and a positive integer $m$.
@@
Moreover the following diagram is commutative, where Res stands for the restriction map.
@@
    $G$ and $H$ are quadratic and ${\mathbb{Q}}(G)={\mathbb{Q}}(H)$.
In that case $\mathcal{R}_{G\times H}=\pi^{-1}_{G\times H,G}(\mathcal{R}_{G})\cap\pi^{-1}_{G\times H,H}(\mathcal{R}_{H})$ and $\mathcal{S}_{G\times H}=\pi^{-1}_{G\times H,G}(\mathcal{S}_{G})\cap\pi^{-1}_{G\times H,H}(\mathcal{S}_{H})$. Moreover, if (2) holds, then $G\times H$ is quadratic and ${\mathbb{Q}}(G\times H)={\mathbb{Q}}(G)$.
@@
Now suppose that $G$ and $H$ are quadratic and ${\mathbb{Q}}(G)={\mathbb{Q}}(H)=K$. By the previous case, we may assume that $K\neq{\mathbb{Q}}$, and we have to show that $G\times H$ is quadratic, $\mathcal{R}_{G\times H}=\pi^{-1}_{G}(\mathcal{R}_{G})=\pi_{H}^{-1}(\mathcal{R}_{H})$ and $\mathcal{S}_{G\times H}=\pi^{-1}_{G}(\mathcal{S}_{G})=\pi^{-1}_{H}(\mathcal{S}_{H})$. By Proposition 3.13, ${\bf Z}({\mathbb{Q}}G)$ is isomorphic to ${\mathbb{Q}}^{n_{1}}\times K^{m_{1}}$ and ${\bf Z}({\mathbb{Q}}H)$ is isomorphic to ${\mathbb{Q}}^{n_{2}}\times K^{m_{2}}$ for some non-negative integers $n_{1},n_{2},m_{1}$ and $m_{2}$. Then ${\bf Z}({\mathbb{Q}}(G\times H))\cong{\bf Z}({\mathbb{Q}}G)\otimes_{\mathbb{Q}}{\bf Z}({\mathbb{Q}}H)\cong{\mathbb{Q}}^{n}\times K^{m}$ where $n=n_{1}n_{2}$ and $m=n_{1}m_{2}+n_{2}m_{1}+2m_{1}m_{2}$. So, by Proposition 3.13, the group $G\times H$ is quadratic and $K={\mathbb{Q}}(G\times H)$. By equations (2.1) and (3.2), we have that $\mathcal{R}_{G\times H}=\sigma_{G\times H}^{-1}(\operatorname{Gal}({\mathbb{Q}}_{G\times H}/K))=\pi^{-1}_{G}(\sigma_{G}^{-1}(\operatorname{Gal}({\mathbb{Q}}_{G}/K)))=\pi^{-1}_{G}(\mathcal{R}_{G})$. Moreover, by Proposition 3.13,
@@
Let $G$ be a USR group which is not quadratic. If $G=G_{1}\times G_{2}$, then either $G_{1}$ or $G_{2}$ is rational.
We conclude the section with the second characterization of quadratic groups.